Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

OG2016 DS A box contains only

Expert replies
by lionsshare » Mon Aug 28, 2017 3:47 pm
A box contains only red chips, white chips, and blue chips. If a chip is randomly selected from the box, what is the probability that the chip will be either white or blue?

(1) The probability that the chip will be blue is 1/5.
(2) The probability that the chip will be red is 1/3.

OA: B

Hi, Experts. Please, anyone, explain the solution for this problem. Thank you.
Join the discussion
Source: — Data Sufficiency |

by ceilidh.erickson » Tue Aug 29, 2017 8:20 am
If the box contains only red, white, and blue chips:
(probability of selecting red) + (probability of selecting white) + (probability of selecting blue) = 1

The questions asks us for the probability of white or blue chips, so we can treat the group (white or blue) as simply (non-red).
(probability of selecting red) + (probability of selecting NON-red) = 1

Thus, if we know the probability of selecting a red, we'll know the probability of selecting anything else (white or blue).

Target question: what is the probability of selecting red?

(1) The probability that the chip will be blue is 1/5.
This gives us blue alone, but not white. Insufficient.

(2) The probability that the chip will be red is 1/3.
Sufficient. This answers our target question. If we know that the probability of selecting red is 1/3, then the probability of anything else (white or blue) must be 2/3.

The answer is B.
Ceilidh Erickson
EdM in Mind, Brain, and Education
Harvard Graduate School of Education
Join the discussion

by Jay@ManhattanReview » Tue Aug 29, 2017 10:30 pm
lionsshare wrote:A box contains only red chips, white chips, and blue chips. If a chip is randomly selected from the box, what is the probability that the chip will be either white or blue?

(1) The probability that the chip will be blue is 1/5.
(2) The probability that the chip will be red is 1/3.

OA: B

Hi, Experts. Please, anyone, explain the solution for this problem. Thank you.
Since the box contains only red, white, and blue chips,

(probability of selecting a red chip) + (probability of selecting a white chip) + (probability of selecting a blue chip) = 1

Say probability of selecting a red chip = r, probability of selecting a white chip = w, probability of selecting a blue chip = b

Thus, r + w + b = 1

We have to find out the probability that the chip will be either white or blue.

Or, we have to find out the value of (w + b).

Statement 1: The probability that the chip will be blue is 1/5.

=> b = 1/5. But we do not know the value of w. Insufficient.

Statement 2: The probability that the chip will be red is 1/3.

=> r = 1/3

=> w + b = 1 – r = 1 – 1/3 = 2/3. Sufficient.

The correct answer: B

Hope this helps!

Download free ebook: Manhattan Review GMAT Quantitative Question Bank Guide

-Jay
_________________
Manhattan Review GMAT Prep

Locations: New York | Barcelona | Manila | Melbourne | and many more...

Schedule your free consultation with an experienced GMAT Prep Advisor! Click here.
Join the discussion

by Matt@VeritasPrep » Thu Aug 31, 2017 5:13 pm
The trick here:

P(W) + P(B) + P(R) = 1

We want P(W) + P(B). Notice that by subtracting P(R) from both sides of the equation above, we have

P(W) + P(B) = 1 - P(R)

So knowing P(R) will give us P(W) + P(B)! Once we're told in S2 that P(R) = 1/3, we've got

P(W) + P(B) = 1 - 1/3

P(W) + P(B) = 2/3

and we're set!
Join the discussion